# MMUTF: Multimodal Multimedia Event Argument Extraction with Unified Template Filling

### v1

@@ 1 Introduction
As the creation of multimodal EE datasets is time-consuming and costly, recent advanced MEE methods focus on (weak) image-text alignment based on available unimodal task-dependent and task-independent vision-language datasets (Liu et al., 2024). However, current multimedia EAE models are still based on simple classification techniques while ignoring cross-ontology transfer capabilities and event template semantics.
@@
We summarize our main contributions as follows: 1) We introduce MMUTF, a unified template filling framework that addresses the EAE task for multiple input modalities. 2) We validate our approach on the widely used M2E2 benchmark and demonstrate superior performance compared to the majority of baselines. 3) We analyze the transfer learning capabilities for multimedia EAE by employing FrameNet as a resource-rich data source.
@@ 2 Method
For event argument roles, the extraction task can be framed as a candidate-query matching problem, where argument roles serve as the queries. We can utilize the matching between candidates and queries to address the multimodal EAE task with a unified template filling model. Specifically, we aim to create candidate representations for either textual entities or image objects and compute a matching score for each query (i.e., argument role) in a given event template. [^1] We can also consider text spans, image segments, etc., as candidates. However, entities and objects follow the present annotation schemes and suit the EAE task best.
@@ 2.1 Multimodal Unified Template Filling
In order to obtain modality-specific candidates, we recognize entities and objects with off-the-shelf systems and extract contextualized representations using a Transformer model for each modality (Vaswani et al., 2017; Raffel et al., 2020; Radford et al., 2021). The encoding process yields a sequence of token embeddings $t=\{t_{i}\}_{1}^{N}\in\mathbb{R}^{N\times H}$ and patch embeddings $p=\{p_{i}\}_{1}^{M}\in\mathbb{R}^{M\times H}$ with dimensionality $H$ for each input sentence $S$ and image $I$, respectively. For entities tokenized into multiple subwords, we use mean pooling and concatenate the feature vector with the representation of the event trigger word. Similarly, for images, we obtain object feature vectors using the max pooling [^2] Max pooling achieved the best results among other techniques such as mean and RoI pooling. operation over all patches containing the detected object, and concatenate these with the image CLS embeddings for further processing. The concatenated vectors are then transformed via their corresponding mapping networks, which will be described in more detail later.
@@
For event template processing, we employ the decoder of a Transformer Encoder-Decoder model (Raffel et al., 2020) as shown in Figure 1. This approach has shown promising results without compromising on the length of the input context, as seen in encoder architectures (Ma et al., 2022). This configuration enables the prompt to interact with the input contexts at the cross-attention layers of every decoder block, resulting in context-dependent query representations. Similar to the generation of candidate representations, we use the mean pooling operation for argument roles that span multiple subwords and feed the features into a subsequent mapping network.
@@
We aim to align the modality-specific candidate and query representations into a unified latent space, based on the candidate and query feature vectors. However, during training, we face the challenge of unifying the rich and diverse representations of the different modalities and Transformer models. The mapping networks are designed to bridge the gap between the latent spaces and transform all candidate and query features into ${H}$-sized representations. Each mapping network shares the same architecture and consists of a two-layer feedforward network with $4\times H$ hidden units, ReLU activation, and is trained with a dropout probability of 40%.
@@ 2.2 Training and Inference
where $\mathcal{L_{\text{BCE}}}$ denotes the binary cross-entropy loss, $y_{ij}$ the ground truth label, $C$ the collection of candidates and $A$ the set of possible argument roles. At inference time, we use a threshold $\tau$ with $\phi(c,r)\geq\tau$ to decide whether candidate $c$ corresponds to argument role $r$.
@@ 3 Experiments
Datasets
We conduct our experiments on the widely used M2E2 multmedia event extraction benchmark (Li et al., 2020) with 8 event types and 15 argument roles. The dataset contains 245 multimedia documents with 6,167 sentences and 1,014 images. It encompasses 1,297 lexical events and 391 visual events, with 192 lexical events and 203 visual events aligning to 309 multimedia events (cf. Figure 2). Since the M2E2 dataset does not provide a training split, we follow previous work (Li et al., 2020; Liu et al., 2022; Du et al., 2023) and employ ACE2005 and SWiG for the training phase. ACE2005 (Walker, Christopher et al., 2006) consists of 33 textual event types and 36 argument roles. SWiG (Pratt et al., 2020) annotates visual events with 504 activity verbs and 1,788 semantic roles. We align both datasets with the M2E2 event ontology.
Metrics
Baselines
We compare our proposed approach with a wide range of state-of-the-art models. These models include FLAT (Li et al., 2020), WASE (Li et al., 2020), CLIPEvent (Li et al., 2022), UniCL (Liu et al., 2022), CAMEL (Du et al., 2023), MGIM (Liu et al., 2024), and UMIE (Sun et al., 2024). The majority of baselines set the focus on weak image-text alignment but neglect more sophisticated classification approaches w.r.t. EAE. We provide a more comprehensive overview in A.2.
Experimental Setup
Our system is trained jointly for 5 epochs on the visual task and 25 epochs on the lexical task with a batch size of 16 and 32, respectively. The different number of epochs reduces the imbalance of modalities in the training datasets. For optimization, we use AdamW (Loshchilov and Hutter, 2019) with an initial learning rate of $3\times 10^{-5}$, weight decay of $1\times 10^{-3}$, and a linear scheduler. We select the best model checkpoint based on F1 of the ACE2005 test dataset and set the default threshold $\tau=0.5$ for inference. We report threshold tuning and implementation details in A.3 and A.4.
3.1 Results and Analysis
@@
Table 1: Comparison with SOTA methods on the M2E2 benchmark for textual, visual, and multimedia events. Bold numbers indicate the best EAE model whereas underlined metrics denote the second best. *UMIE is not directly comparable due to a lack of evaluation details.
@@
The results of our main experiments can be found in Table 1. In the following sections, we present observations and further analysis.
Effectiveness of MMUTF
We find that our model surpasses the SOTA in terms of textual EAE by +7% F1. For visual and multimedia EAE, we generally perform better than the second-best systems, achieving improvements of +3.1% and +2.8% F1, respectively. These results demonstrate the efficacy of our approach by utilizing cross-ontology transfer capabilities and event semantics via natural language prompts. However, we observe a gap of -5.8% F1 in multimedia EAE compared to CAMEL, which is mainly caused by the visual modality. We hypothesize that CAMEL’s better performance is related to its additional data augmentation pipeline and the integration of textual clues. Note that we do not utilize any additional image-language paired datasets or cross-modal data augmentation in this work.
@@
Table 2 presents results for different multimodal training strategies, focusing on cross-modality alignment for EAE. Text + Image denotes the training strategy that jointly trains the lexical and visual input modalities. Text $\rightarrow$ Image represents the sequential procedure. Here, we first train our model on the text input modality, then freeze the text and query model parameters before training the visual input modality. With Image $\rightarrow$ Text, we follow the same procedure as before, but in reverse. Lastly, we follow recommendations in vision-language alignment (Zhai et al., 2022) and freeze the visual encoder during joint training (Image Locked). The results suggest that joint training achieves superior performance, while the sequential procedure experiences performance drops of up to -1.9% for the aligned modality. The Image Locked approach yields the best textual and multimodal F1 scores, while also representing the most memory-efficient alternative during training.
@@
Table 3: Evaluaton of FrameNet transfer learning.
Transfer Learning from SRL
Previous work has shown that semantic role labeling (SRL) can serve as a valuable resource for EAE (Zhang et al., 2022). Analogous task similarities are also present in image activity recognition, which utilizes FrameNet verbs and frame elements as semantic roles (Yatskar et al., 2016). However, no work has yet explored the transfer learning capabilities for multimedia EAE. We use FrameNet 1.7 (Baker et al., 1998) and SWiG (Pratt et al., 2020) as our training datasets and merge their ontologies to 155 frames. Due to the large number of frames, the event templates are constructed with a semi-automatic approach (Zhang et al., 2022) (cf. A.5). Table 3 shows the results for zero-shot prediction on the M2E2 benchmark with gold and predicted event triggers. We observe remarkable transfer learning capabilities without any training on the M2E2 ontology. In fact, this simplistic approach outperforms even the supervised baselines FLAT and WASE by up to +11.6% and +5.8% F1 for visual and multimedia EAE, respectively.
@@ 4 Conclusion
In this work, we investigated the multimedia EAE task and proposed MMUTF, a simple yet effective model that assigns candidates to corresponding argument roles. Our results demonstrated the effectiveness of the model and highlight the benefits of cross-ontology transfer and event semantic incorporation. This assumption was supported by SRL transfer learning experiments, which we adapted to the multimedia EAE task. Future work will focus on incorporating additional textual and visual clues from input documents, as well as exploring bi-encoder architectures to reduce inference time.
@@
In this work, we propose a unified template-filling approach for solving multimodal multimedia event argument extraction. Our method relies on manually crafted prompts that represent the event templates for each event type. The time required to construct such prompts was reasonable for the M2E2 benchmark, which covers only 8 event types and 15 argument roles for solving this task. Other event extraction ontologies cover over 100 event types and complex argument relations (Huang et al., 2023). Although we used a semi-automatic approach for the more comprehensive FrameNet, we still identified erroneous event templates, indicating that this method is not fully reliable. The manual prompt construction for such large event ontologies also poses challenges in terms of required time and complexity. In addition, our system relies on event predictions, textual entities, and visual objects generated by preceding systems. This approach is prone to error propagation and requires a holistic assessment of all system components with appropriate evaluation data.
@@ Appendix A Appendix
A.1 Event Example
In Figure 2, we depict a representative example from the M2E2 benchmark (Li et al., 2020). Event argument roles such as vehicle and destination are only present in the visual and textual modality, respectively.
Figure 2: An example of Multimedia Event Extraction from the M2E2 benchmark, where event argument roles are extracted from both the textual and visual modality.
A.2 Baselines
We compare our proposed approach with a wide range of SOTA models. 1) WASE (Li et al., 2020) uses graph neural networks (GNN) together with abstract meaning representations and situation graphs for text and image, respectively. These models additionally include the VOA image-caption dataset for weak cross-modality alignment, while the 2) FLAT (Li et al., 2020) baseline removes the GNN components. 3) CLIPEvent (Li et al., 2022) combines the previous approaches with pre-trained CLIP model weights and introduces a Optimal Transport strategy for event structure representation and modality alignment. 4) UniCL (Liu et al., 2022) proposes a unified contrastive learning framework and uses a contrastive learning mechanism and incorporates textual and visual clues. 5) CAMEL (Du et al., 2023) is based on cross-modal data augmentation, integrates an adapter module and applies a specialized training strategy. 6) MGIM (Liu et al., 2024) is a multi-grained gradual inference model with fine-grained alignment between texts and images by multiple rounds of gradual inference. 7) UMIE (Sun et al., 2024) utilizes instruction tuning for unified multimodal information extraction based on text generation and Flan-T5 in various model sizes.
A.3 Threshold Tuning
Figure 3: EAE F1 results for our MUTF model with varying thresholds for each modality.
We investigate the influence of an automatic or manually chosen threshold for both the textual and visual matching scores. In our work, we select the default threshold $\tau=0.5$ and obtain reasonable results. Figure 3 shows the metrics for different thresholds and modalities. Additionally, we analyze threshold tuning based on the best F1-scores for the ACE and SWiG test datasets. This procedure yields $\tau^{T}=0.2$ and $\tau^{V}=0.3$ for text and vision, respectively. For automatic threshold tuning, we obtain a multimedia F1-score of 28.3%, improving our default $\tau$ by +0.9%. In addition, we find the most performance degradation above a threshold of 0.6 while achieving the best results between 0.3 and 0.5. This suggests a weak discriminative power between positive and negative candidate-query pairs. Contrastive learning and hard negative mining may address this issue, but it is not within the scope of this work.

### v2

@@ 1 Introduction
Figure 1: Example of Multimedia Event Extraction from the M2E2 benchmark, where event argument roles are extracted from both the textual and visual modality.
@@
As the creation of multimodal EE datasets is time-consuming and costly, recent advanced MEE methods focus on (weak) image-text alignment based on available unimodal task-dependent and task-independent vision-language datasets (Li et al., 2020; Liu et al., 2022; Du et al., 2023; Liu et al., 2024). However, current multimedia EAE models are still based on simple classification techniques while ignoring cross-ontology transfer capabilities and event template semantics.
@@
We summarize our main contributions as follows: 1) We introduce MMUTF, a unified template filling framework that addresses the EAE task for multiple input modalities. 2) We validate our approach on the widely used M2E2 benchmark and demonstrate superior performance compared to the majority of baselines. 3) We analyse the transfer learning capabilities for multimedia EAE by employing FrameNet as a resource-rich data source.
@@ 2 Method
For event argument roles, the extraction task can be framed as a candidate-query matching problem, where argument roles serve as the queries. We can utilize the matching between candidates and queries to address the multimedia EAE task with a unified template filling model. Specifically, we aim to create candidate representations for either textual entities or image objects and compute a matching score for each query (i.e., argument role) in a given event template. [^1] We can also consider text spans, image segments, etc., as candidates. However, entities and objects follow the present annotation schemes and suit the EAE task best.
@@ 2.1 Multimodal Unified Template Filling
In order to obtain modality-specific candidates, we recognize entities and objects with off-the-shelf systems and extract contextualized representations using a Transformer model for each modality (Vaswani et al., 2017; Raffel et al., 2020; Radford et al., 2021). The encoding process yields a sequence of $N$ token embeddings $t=\{t_{i}\}_{1}^{N}\in\mathbb{R}^{N\times H}$ and $M$ patch embeddings $p=\{p_{i}\}_{1}^{M}\in\mathbb{R}^{M\times H}$ with dimensionality $H$ for each input sentence $S$ and image $I$, respectively. For entities tokenized into multiple subwords, we use mean pooling and concatenate the feature vector with the representation of the event trigger word. Similarly, for images, we obtain object feature vectors using the max pooling [^2] Max pooling achieved the best results among other techniques such as mean and RoI pooling. operation over all patches containing the detected object, and concatenate these with the image CLS embeddings for further processing. Here, the CLS embeddings serve as event triggers represented as image scenes. The concatenated vectors are then transformed via their corresponding mapping networks, which will be described in more detail later.
@@
For event template processing, we employ the decoder of a Transformer Encoder-Decoder model (Raffel et al., 2020) as shown in Figure 2. This approach has shown promising results without compromising on the length of the input context, as seen in encoder architectures (Ma et al., 2022). This configuration enables the prompt to interact with the input contexts at the cross-attention layers of every decoder block, resulting in context-dependent query representations. Similar to the generation of candidate representations, we use the mean pooling operation for argument roles that span multiple subwords and feed the features into a subsequent mapping network. We denote the query representations as $q=\{q_{i}\}_{1}^{R}\in\mathbb{R}^{R\times H}$, where $H$ is the dimensionality, and $R$ is the number of argument roles.
@@
We aim to align the modality-specific candidate and query representations into a unified latent space, based on the candidate and query feature vectors. However, during training, we face the challenge of unifying the rich and diverse representations of the different modalities and Transformer models. The mapping networks are designed to bridge the gap between the latent spaces and transform all candidate and query features into ${H}$-sized representations. Each mapping network shares the same architecture and consists of a two-layer feedforward network with $4\times H$ hidden units, ReLU (Agarap, 2018) activation, and is trained with a dropout probability of 40%.
@@ 2.2 Training and Inference
where $\mathcal{L_{\text{BCE}}}$ denotes the binary cross-entropy loss, $y_{ij}$ the ground truth label, $C$ the collection of candidates and $A$ the set of possible argument roles. At inference time, we use a threshold $\tau$ with $\phi(c,r)\geq\tau$ to decide whether candidate $c$ corresponds to argument role $r$. For each event and candidate, we assign the role with the highest score that exceeds the threshold $\tau$. Otherwise, the candidate is assigned to the None class.
@@ 3 Experiments
Table 1: Comparison with SOTA methods on the M2E2 benchmark for textual, visual, and multimedia events. Bold numbers indicate the best EAE model whereas underlined metrics denote the second best. $\dagger$ denotes our reproduced results and * is not directly comparable due to lack of evaluation details.
Datasets
We conduct our experiments on the widely used M2E2 multmedia event extraction benchmark (Li et al., 2020) with 8 event types and 15 argument roles. The dataset contains 245 multimedia documents with 6,167 sentences and 1,014 images. It encompasses 1,297 textual events and 391 visual events, with 192 textual events and 203 visual events aligning to 309 multimedia events (cf. Figure 1). Since the M2E2 dataset does not provide a training split, we follow previous work (Li et al., 2020; Liu et al., 2022; Du et al., 2023) and employ ACE2005 and SWiG for the training phase. ACE2005 (Walker, Christopher et al., 2006) consists of 33 textual event types and 36 argument roles. SWiG (Pratt et al., 2020) annotates visual events with 504 activity verbs and 1,788 semantic roles. Following Li et al. (2020), we align both datasets with the M2E2 event ontology.
Metrics
Baselines
We compare our proposed approach with a wide range of state-of-the-art models. These models include FLAT (Li et al., 2020), WASE (Li et al., 2020), CLIPEvent (Li et al., 2022), UniCL (Liu et al., 2022), CAMEL (Du et al., 2023), MGIM (Liu et al., 2024), and UMIE (Sun et al., 2024). The majority of baselines set the focus on weak image-text alignment but neglect more sophisticated classification approaches w.r.t. EAE. We provide a more comprehensive overview in A.1.
Experimental Setup
Our system is trained jointly for 5 epochs on the visual task and 25 epochs on the textual task with a batch size of 16 and 32, respectively. The varying number of epochs reduces the imbalance of modalities in the training datasets. For optimization, we use AdamW (Loshchilov and Hutter, 2019) with an initial learning rate of $3\times 10^{-5}$, weight decay of $1\times 10^{-3}$, and a linear scheduler. We select the best model checkpoint based on F1 of the ACE2005 test dataset and set the default threshold $\tau=0.5$ for inference. We report threshold tuning and implementation details in A.2 and A.3.
3.1 Results
@@
The results of our main experiments can be found in Table 1. We find that our model surpasses the SOTA in terms of textual EAE by +7% F1. For visual and multimedia EAE, we generally perform better than the second-best systems, achieving improvements of +3.1% and +2.8% F1, respectively. These results demonstrate the efficacy of our approach by utilizing cross-ontology transfer capabilities and event semantics via natural language prompts. However, we observe a gap of -5.8% F1 in multimedia EAE compared to CAMEL, primarily due to the visual modality. With the same set of predicted event mentions, the gap narrows to -0.5% F1 in multimedia EAE. We hypothesize that CAMEL’s better visual performance is related to its additional data augmentation pipeline and the integration of textual clues. Note that we do not utilize any additional image-language paired datasets or cross-modal data augmentation in this work. Incorporating cross-modal information from the document’s context and external datasets, as well as enhanced event detection, might further improve our model’s performance.
3.2 Analysis
@@
Table 2 presents results for different multimodal training strategies, focusing on cross-modality alignment for EAE. Text + Image denotes the training strategy that jointly trains the textual and visual input modalities. Text $\rightarrow$ Image represents the sequential procedure. Here, we first train our model on the text input modality, then freeze the text and query model parameters before training the visual input modality. With Image $\rightarrow$ Text, we follow the same procedure as before, but in reverse. Lastly, we follow recommendations in vision-language alignment (Zhai et al., 2022) and freeze the visual encoder during joint training (Image Locked). The results suggest that joint training achieves superior performance, while the sequential procedure experiences performance drops of up to -1.9% F1 for the aligned modality. The Image Locked approach yields the best textual and multimodal F1 scores, while also representing the most memory-efficient alternative during training.
Ablation Study
To better illustrate the effectiveness of different components, we conduct ablation studies and present the results in Table 3. In line 1, we remove the cross-attention mechanism between the candidate and query representations. This results in a performance reduction for visual and multimedia EAE, while we observe a slight increase for textual EAE. In line 2, we replace the joint model for query representations with two modality-specific models. The performance of all EAE cases significantly declines by up to +3.4% in terms of F1 scores. These observations suggest the effectiveness of unified template filling, particularly for the visual and multimedia EAE tasks. Lastly, to assess the influence of natural language-formulated event templates, we replace the textual query representations with trainable argument role prototypes in line 3. The performance declines mostly for visual and multimedia EAE, indicating that event template semantics provide beneficial information to enhance cross-ontology transfer capabilities. We provide backbone and prompt template ablations in A.5 and A.6, respectively.
Model Textual F1 Visual F1 MM F1
MMUTF 38.2 20.9 27.4
1   w/o Cross-Attention 38.5 17.8 26.5
2   w/o Joint-Prompts 37.0 17.5 24.4
3   w/o Prompts 37.9 15.9 23.3
Table 3: Ablations for the different components. ($1$) without cross-attention between candidate and query representations. ($2$) Replacement of joint query model with modality-specific models. ($3$) Replacement of textual query representations with trainable prototypes.
Transfer Learning from SRL
Previous work has shown that semantic role labeling (SRL) can serve as a valuable resource for EAE (Zhang et al., 2022). Analogous task similarities are also present in image activity recognition, which utilizes FrameNet verbs and frame elements as semantic roles (Yatskar et al., 2016). However, no work has yet explored the transfer learning capabilities for multimedia EAE. We use FrameNet 1.7 (Baker et al., 1998) and SWiG (Pratt et al., 2020) as our training datasets and merge their ontologies to 155 frames. Due to the large number of frames, the event templates are constructed with a semi-automatic approach (Zhang et al., 2022) (cf. A.4). Table 4 shows the results for zero-shot prediction on the M2E2 benchmark with gold and predicted event triggers. We observe remarkable transfer learning capabilities without any training on the M2E2 ontology. In fact, this simplistic approach outperforms even the supervised baselines FLAT and WASE by up to +11.6% and +5.8% F1 for visual and multimedia EAE, respectively.
@@
Table 4: Evaluaton of FrameNet transfer learning. Gold Triggers represent EAE based on annotated event mentions. Pred Triggers represent EAE based on predicted event mentions.
@@ 4 Conclusion
In this work, we investigate the multimedia EAE task and propose MMUTF, a simple yet effective unified template filling approach that assigns candidates to corresponding argument roles. We encode entity candidates and object candidates from textual and visual inputs, respectively. Then, we feed a natural language event template into the query decoder, which facilitates interactions with document contexts via cross-attention mechanisms. Finally, candidate and query representations serve as the basis for computing candidate-query matching scores. Our results demonstrate the effectiveness of the model and highlight the benefits of cross-ontology transfer and event semantic incorporation. This assumption is supported by SRL transfer learning experiments, which we adapt to the multimedia EAE task. Future work will focus on incorporating additional textual and visual cues from input documents, integrating event detection, and exploring bi-encoder architectures to reduce inference time.
@@
In this work, we propose a unified template-filling approach for solving multimodal multimedia event argument extraction. Our method relies on manually crafted prompts that represent the event templates for each event type. The time required to construct such prompts was reasonable for the M2E2 benchmark, which covers only 8 event types and 15 argument roles for solving this task. Other event extraction ontologies cover over 100 event types and complex argument relations (Huang et al., 2023). Although we used a semi-automatic approach for the more comprehensive FrameNet, we still identified erroneous event templates, indicating that this method is not fully reliable. The manual prompt construction for such large event ontologies also poses challenges in terms of required time and complexity. In addition, our system relies on event predictions, textual entities, and visual objects generated by preceding systems. This approach is prone to error propagation (cf. A.7) and requires a holistic assessment of all system components with appropriate evaluation data.
@@ Appendix A Appendix
A.1 Baselines
We compare our proposed approach with a wide range of SOTA models. 1) WASE (Li et al., 2020) uses graph neural networks (GNN) together with abstract meaning representations and situation graphs for text and image, respectively. These models additionally include the VOA image-caption dataset for weak cross-modality alignment, while the 2) FLAT (Li et al., 2020) baseline removes the GNN components. 3) CLIPEvent (Li et al., 2022) combines the previous approaches with pre-trained CLIP model weights and introduces a Optimal Transport strategy for event structure representation and modality alignment. 4) UniCL (Liu et al., 2022) proposes a unified contrastive learning framework, uses a multimodal contrastive learning mechanism and incorporates textual and visual clues. 5) CAMEL (Du et al., 2023) is based on cross-modal data augmentation, integrates an adapter module and applies a specialized training strategy. 6) MGIM (Liu et al., 2024) is a multi-grained gradual inference model with fine-grained alignment between texts and images by multiple rounds of gradual inference. 7) UMIE (Sun et al., 2024) utilizes instruction tuning for unified multimodal information extraction based on text generation and Flan-T5 (Chung et al., 2024) in various model sizes.
A.2 Threshold Tuning
We investigate the influence of an automatic or manually chosen threshold for both the textual and visual matching scores. In our work, we select the default threshold $\tau=0.5$ and obtain reasonable results. Figure 3 shows the metrics for different thresholds and modalities. Additionally, we analyze threshold tuning based on the best F1-scores for the ACE and SWiG test datasets. This procedure yields $\tau^{T}=0.2$ and $\tau^{V}=0.3$ for text and vision, respectively. For automatic threshold tuning, we obtain a multimedia F1-score of 28.3%, improving our default $\tau$ by +0.9%. In addition, we find the most performance degradation above a threshold of 0.6 while achieving the best results between 0.3 and 0.5. This suggests a weak discriminative power between positive and negative candidate-query pairs. Contrastive learning and hard negative mining may address this issue, but is not within the scope of this work.
Figure 3: EAE F1 results for our MMUTF model with varying thresholds for each modality.